OpenStudy (anonymous):

What is the radius of a circle with a circumference of twenty-six point six nine meters? Use three point one four for pi and show your work.

OpenStudy (lucaz):

\[\frac{Circunference}{diameter}=\pi\] diameter = (2)radius \[\frac{26.69}{2\times radius}=3.14\] solve for radius..